On the basis of 5 years of data the following probability function for a company’s

 

weekly demand for wool was:

Amount of wool (kg) 2500 3500 4500 5500

 Probability 0.35 0.45 0.20 0.05

What was the expected weekly demand for wool based on the distribution?


1
Expert's answer
2021-04-29T16:58:14-0400

he expected weekly demand for wool is equal to the expected value:

"M(x) = \sum {{x_i}} {p_i} = 2500 \cdot 0.35 + 3500 \cdot 0.45 + 4500 \cdot 0.2 + 5500 \cdot 0.05 = 3625"

Answer: 3625 kg